# Why Rectified Power Unit Networks Fail and How to Improve It: An Effective Theory Perspective

### v3

@@ 5 Approximation Properties of MRePU networks
In this section, we theoretically demonstrate that a neural network constructed with the MRePU activation function can reconstruct polynomials exactly, similar to RePU neural networks, and achieves the same universal approximation property for the class of differentiable functions as shown in [15]. Additionally, we experimentally verify that, beyond this theoretical foundation, neural networks with the MRePU activation function possess an appropriate inductive bias when approximating certain classes of functions.
@@ 5.1 Approximation Theorems for MRePU
In [15], it is shown that a RePU network of $p$-th order can accurately represent an expanded univariate polynomial using Horner’s method, utilizing $4p$ nodes for multiplication and $2p$ nodes to construct a $p$-th order polynomial. Furthermore, by applying these univariate polynomial results inductively, the representation of multivariate polynomials can also be achieved. Therefore, for the MRePU of order $p$ network, if we can demonstrate that it can construct any $p$-th order polynomial using $2p$ nodes, as in the case of RePU, we obtain the same result. The following lemma formalizes this.
Lemma 7 A polynomial of degree $p$, $\sum_{k=0}^{p}a_{k}x^{k}$ can be approximated by $p$-th order MRePU network with one hidden layer and nodes no more than $2p$.
@@
In this case, since the components of $W_{0}^{T}x+b_{0}$ where the MRePU activation is applied are symmetric with respect to $x=-1$, only one set of components—either the even-indexed or odd-indexed—remains active. If the odd components remain active, the final output takes the following form:
$v_{k}(x+u_{k})^{p}(x+u_{k}-1).$
In the case of the even components, the output is as follows:
$v_{k}(x+u_{k})^{p}(-x-u_{k}-1).$
In both cases, after binomial expansion, the expression can be written as follows, where $s$ is a value determined by the parity of the index:
$v_{k}\Big{(}(-1)^{s}\sum_{j=0}^{p+1}\binom{p+1}{j}u_{k}^{p+1-j}x^{j}-\sum_{j=0}^{p}\binom{p}{j}u_{k}^{p-j}x^{j}\Big{)}.$
Then, by comparing the final expression of the neural network with the coefficients of the polynomial we aim to construct, $\sum_{k=0}^{p}a_{k}x^{k}$, for $k<p+1$, we obtain the following result:
$\Big{(}(-1)^{s}\binom{p+1}{k}-\binom{p}{k}\Big{)}\sum_{l=1}^{p}(u_{l}^{p+1-k}-u_{l}^{p-k})v_{l}=a_{k}.$
This can be expressed as a matrix equation as follows:
$\begin{bmatrix}u_{1}-1&\dots&u_{p}-1&0\\ \vdots&&\vdots&\vdots\\ u_{1}^{p}-u_{1}^{p-1}&\dots&u_{p}^{p}-u_{p}^{p-1}&0\\ u_{1}^{p+1}-u_{1}^{p}&\dots&u_{p}^{p+1}-u_{p}^{p-1}&1\par\end{bmatrix}\begin{bmatrix}v_{1}\\ \vdots\\ v_{p}\\ v_{0}\end{bmatrix}=\begin{bmatrix}a_{p}\Big{(}(-1)^{s_{p}}\binom{p+1}{p}-\binom{p}{p}\Big{)}^{-1}\\ \vdots\\ a_{1}\Big{(}(-1)^{s_{1}}\binom{p+1}{1}-\binom{p}{1}\Big{)}^{-1}\\ a_{0}\Big{(}(-1)^{s_{0}}\binom{p+1}{0}-\binom{p}{0}\Big{)}^{-1}\end{bmatrix}.$ (9)
Then, the determinant of the top-left $p\times p$ sub-matrix of this equation is given by:
$\prod_{k=1}^{p}(u_{k}-1)\det\bigg{(}\begin{bmatrix}1&\dots&1\\ u_{1}&&u_{p}\\ \vdots&&\vdots\\ u_{1}^{p-1}&\dots&u_{p}^{p-1}\end{bmatrix}\bigg{)}$
The arguments of the determinant on the right form a Vandermonde matrix, so equation (9) has a solution if each $u_{i}$ is distinct and not equal to 1. To make the $p+1$-th order term vanish, it suffices to show the coefficient $(-1)^{s_{1}}v_{1}+\dots+(-1)^{s_{p}}v_{p}$ can be reduced. It can be easily verified that the hyperdeterminant of the Jacobian of the top-left $k\times k$ sub-matrix of equation (9) is a constant multiple of the determinant of the Vandermonde matrix. Therefore, by the inverse function theorem, we can conclude that the following function occupies a positive measure in a neighborhood containing the origin in $\mathbb{R}^{p}$ space.
$F(u_{1},\dots,u_{p})=\begin{bmatrix}u_{1}-1&\dots&u_{p}-1\\ \vdots&&\vdots\\ u_{1}^{p}-u_{1}^{p-1}&\dots&u_{p}^{p}-u_{p}^{p-1}\par\end{bmatrix}^{-1}\begin{bmatrix}a_{p}\Big{(}(-1)^{s_{p}}\binom{p+1}{p}-\binom{p}{p}\Big{)}^{-1}\\ \vdots\\ a_{1}\Big{(}(-1)^{s_{1}}\binom{p+1}{1}-\binom{p}{1}\Big{)}^{-1}\end{bmatrix}.$
Therefore, by choosing sufficiently small $v_{i}$’s, the coefficient of the $p+1$-th order monomial can be reduced to the desired degree. $\square$
@@
Then, by following this lemma and the proof outlined in [15], we obtain the following theorem:
@@
Theorem 8 if $f:\mathbb{R}^{d}\rightarrow\mathbb{R}$ is a polynomial of $d$ variables with degree $N$, then $f$ can be exactly reconstructed via MRePU neural network with $12N-$ 1 hidden layers, $(6p+2)(2N^{d}-N^{d-1}-N)+2p(2N^{d}-N^{d-1}-N)/(N-1)\simeq\mathcal{O}(pN^{d})$ number of neurons, $(30p+2)(2N^{d}-N^{d-1}-N)+(2p+1)(2N^{d}-N^{d-1}-N)/(N-1)\simeq\mathcal{O}(pN^{d})$ number of parameters and width $12pN^{d-1}+6p(N^{d-1}-N)/(N-1)\simeq\mathcal{O}(pN^{d-1})$.
@@
Additionally, as a corollary to the above theorem, we obtain a universal approximation theorem for the class of differentiable functions. Before explaining this, we define a reference norm that will serve as the standard. Definition 9 We define the norm of $f\in C^{s}(A)$, where $A\subset\mathbb{R}^{d}$, as following:
@@
Corollary 10 Let $f$ be a function belongs to $C^{s}(K)$ where $K\subset\mathbb{R}^{d}$ is a compact set. For any $N\in\mathbb{N}^{+}$, there exists a MRePU neural network $\psi_{N}$ width depth $\mathcal{D}$, width $\mathcal{W}$, number of neurons $\mathcal{U}$ and total number of parameters $\mathcal{P}$ with following specification:
@@ 5.2 Experiments
Regression on polynomials and differentiable functions First, we validate the efficiency of MRePU on polynomial functions. The neural network architecture consists of five hidden layers, each with a width of 32, and the input dimension is set according to the variables in each polynomial. For univariate polynomial problems, the training dataset was constructed with inputs from -2.0 to 2.0, divided into 101 points at intervals of 0.04. For $x^{3}-x$, the learning rate was set to 0.005, and $C_{W}$ was set to 0.5 for MRePU and 1.0 for other cases. For $x^{5}-5x^{3}+6x$, the learning rate was set to 0.0001, with $C_{W}$ set to 0.8 for MRePU case and learning rate to 0.001, $C_{W}$ to 1.0 for the others. For multivariate cases, the dataset for $xy$ was created by forming a meshgrid over the range $[-2,2]\times[-2,2]$, with each axis divided into 101 points. Similarly, for $xyz$, a meshgrid was formed over the range $[-1,1]\times[-1,1]\times[-1,1]$, with each axis divided into 51 points. Training was conducted up to 10,000 epochs for all cases. The learning rate for multivariate cases was set to 0.001, and $C_{W}$ was set to 0.55 for $xy$, 0.6 for $xyz$, and 1.0 for the other cases. Each hyperparameter was empirically determined to achieve optimal performance.
$x^{3}-x$ $x^{5}-5x^{3}+6x$ $xy$ $xyz$
MRePU $\bf{4.0*10^{-4}}$ $\bf{2.0*10^{-4}}$ $\bf{2.7*10^{-5}}$ $\bf{2.5*10^{-4}}$
ReLU $8.0*10^{-4}$ $7.0*10^{-4}$ $4.0*10^{-4}$ $4.2*10^{-2}$
GELU $7.2*10^{-2}$ $2.1*10^{-3}$ $4.0*10^{-4}$ $4.2*10^{-2}$
@@
Table 4 presents the average final loss over 25 experiments, where neural networks were trained to regress each polynomial function. As observed in the table, MRePU-based neural networks consistently outperform the others in all cases. Notably, for approximating $xyz$, other networks fail to learn effectively, while only the MRePU network achieves meaningful learning. The loss plots for $x^{5}-5x^{3}+6x$ and $xy$ are depicted in Figure 9. Figure 10 illustrates the regression results of $x^{5}-5x^{3}+6x$ using ReLU and MRePU. A key observation is that the trained MRePU network approximates not only the function itself but also its derivative, whereas the ReLU network produces a piecewise function with noticeable discontinuities. Similarly, Figure 11 shows the contour graphs of trained networks with different activation functions for $xy$. The MRePU network demonstrates a highly accurate approximation, while the ReLU network exhibits significant instability in the contours, and GELU shows reduced accuracy.
Figure 9: The training loss plots for each neural network over epochs for the cases of $x^{5}-5x^{3}+6x$ (Left) and $xy$ (Right).
Figure 10: Top: The results of training for regressing $x^{5}-5x^{3}+6x$, where the left plot shows the output of the MRePU with $p=2$ network, and the right plot shows the output of the ReLU network. Bottom: The comparison between the derivative of the trained networks and the actual derivative function, $65x^{4}-15x^{2}+$ 6.
Figure 11: Contour plots for the reference function and the trained neural networks. Top left: MRePU with $p=2$ network. Top right: Reference ($xy$). Bottom left: GELU network. Bottom right: ReLU network.
@@
The training dataset for univariate functions was constructed in the same manner as for the polynomials. For multivariate functions, a meshgrid was generated by dividing the domain $[-1,1]\times[-1,1]$ into 51 intervals along each axis. The learning rate and $C_{W}$ values for these experiments were as follows: For $\lfloor x\rfloor_{+}^{3}$, the MRePU network used a learning rate of 0.0004 and $C_{W}=0.6$, while the other networks used 0.001 and $C_{W}=1.0$. For $f(x)$, The MRePU network employed a learning rate of 0.001 and $C_{W}=0.6$, while the other networks used 0.001 and $C_{W}=1.0$. For $g(x,y)$, the MRePU network used a learning rate of 0.002 and $C_{W}=0.8$, whereas the remaining networks used 0.003 and $C_{W}=1.0$. Each hyperparameter was empirically determined to achieve optimal performance.
@@
MRePU $8.1*10^{-5}$ $7.9*10^{-5}$ $4.0*10^{-4}$
ReLU $\bf{3.2*10^{-5}}$ $\bf{5.6*10^{-5}}$ $\bf{2.0*10^{-4}}$
GELU $2.4*10^{-3}$ $8.0*10^{-4}$ $4.4*10^{-1}$
@@
As shown in Table 5 and Figure 12, for differentiable functions with discontinuities in higher-order derivatives, the ReLU network generally demonstrates the best performance. However, as observed in Figures 13 and 14, even though the training loss of the ReLU network is lower (i.e., its $L^{2}$ norm is smaller), the MRePU network exhibits more stable function approximation in terms of contour smoothness and the behavior of the differentiated function. For the GELU network, it is evident that learning fails entirely for $g(x,y)$.
@@
Case of Physics-Informed Neural Networks (PINNs) Next, we compare and contrast neural networks constructed with different activation functions using the physics-informed neural networks (PINNs) [26] technique, which solves PDE problems by incorporating the PDE itself into the loss function. The PDE used in this experiment is the Burgers equation, given as:
@@
We defined the problem domain as $(x,t)\in[-1,1]\times[0,1]$ with $\nu=\frac{0.01}{\pi}$. The initial condition (the function value at $t=0$) was set to $-\sin{\pi x}$, and the boundary condition was defined as $u(-1,t)=u(1,t)=0$. The PINN loss comprises the $L^{2}$ loss for the boundary condition and the loss for the PDE. The number of points used for loss calculations was as follows: 10,000 randomly sampled collocation points within the domain and 100 points for the boundary condition. The hyperparameter settings included the use of the Adam optimizer with a learning rate of 0.001, betas set to $(0.9,0.999)$, and weight decay set to 0. For $C_{W}$, the value was set to 0.45 for MRePU with $p=2$ and 1.0 for all other cases. The neural network architecture was consistent across experiments, with an input dimension of 2, seven hidden layers, and a width of 64 per layer.
@@
The results of the experiment are presented in Figure 15, 16, and 17. Interestingly, unlike the previous cases with polynomials and differentiable functions, the GELU activation exhibited the best performance, followed by the MRePU network. The ReLU network, however, failed to achieve meaningful learning. Additional experiments with the Tanh activation showed a performance similar to that of the GELU network. From this experiment, it can be concluded that the MRePU network is capable of achieving meaningful learning even under the highly complex training dynamics of the PINN loss [27], [28] and with a deep architecture consisting of 7 hidden layers (9 layers in total).
@@
Polynomial: MRePU $>$ ReLU $>$ GELU
Differentiable function: ReLU $>$ MRePU∗∗ $>$ GELU
PINN for Burgers equation: GELU $>$ MRePU $>$ ReLU

### v4

@@ 5 Approximation Properties of MRePU networks
In this section, we theoretically demonstrate that a neural network constructed with the MRePU activation function can reconstruct polynomials approximately, similar to RePU neural networks, and achieves the same universal approximation property for the class of differentiable functions as shown in (Shen23). Additionally, we experimentally verify that, beyond this theoretical foundation, neural networks with the MRePU activation function possess an appropriate inductive bias when approximating certain classes of functions.
@@ 5.1 Approximation Theorems for MRePU
In (Shen23), it is shown that a RePU network of $p$-th order can accurately represent an expanded univariate polynomial using Horner’s method, utilizing $4p$ nodes for multiplication and $2p$ nodes to construct a $p$-th order polynomial. Furthermore, by applying these univariate polynomial results inductively, the representation of multivariate polynomials can also be achieved. Therefore, for the MRePU of order $p$ network, if we can demonstrate that it can approximate any $p$-th order polynomial using $2p$ nodes, as in the case of RePU, we obtain the same result. The following lemma formalizes this.
Lemma 10. A polynomial of degree $p$, $\sum_{k=0}^{p}a_{k}x^{k}$ can be represented by a $p$-th order MRePU network with one hidden layer and at most $2p$ hidden nodes.
@@
where $u_{1},\dots,u_{p}$ are pairwise distinct real numbers and $v_{0},v_{1},\dots,v_{p}$ are trainable weights. For each pair $(x+u_{k}-1)$ and $(-x-u_{k}-1)$ passed through the $p$-th order MRePU,
$\begin{split}&v_{k}(x+u_{k})^{p}(x+u_{k}-1)+v_{k}(x+u_{k})^{p}(-x-u_{k}-1)\\ &=-2v_{k}(x+u_{k})^{p}.\end{split}$
Summing over $k$ gives
$\begin{split}y(x)&=-2\sum_{k=1}^{p}v_{k}(x+u_{k})^{p}+v_{0}\\ &=-2\sum_{k=1}^{p}\sum_{j=0}^{p}\binom{p}{j}u_{k}^{p-j}x^{j}+v_{0}.\end{split}$
Comparing coefficients with $\sum_{j=0}^{p}a_{j}x^{j}$ yields, for $j=1,\dots,p$,
$2\sum_{l=1}^{p}\binom{p}{k}u_{l}^{p-k}v_{l}=a_{k}.$
This can be written as the linear system:
$\begin{bmatrix}1&\dots&1\\ \vdots&&\vdots\\ u_{1}^{p-2}&\dots&u_{p}^{p-2}\\ u_{1}^{p-1}&\dots&u_{p}^{p-1}\par\end{bmatrix}\begin{bmatrix}v_{1}\\ \vdots\\ v_{p-1}\\ v_{p}\end{bmatrix}=-\frac{1}{2}\begin{bmatrix}a_{p}\binom{p}{p}^{-1}\\ \vdots\\ a_{2}\binom{p}{2}^{-1}\\ a_{1}\binom{p}{1}^{-1}\end{bmatrix}$. (12)
Since $u_{1},\dots,u_{p}$ are distinct, $V(u)$ is invertible, and hence $v_{1},\dots,v_{p}$ are uniquely determined. Finally, matching the constant term $v_{0}=a_{0}+2\sum_{k=1}^{p}v_{k}u_{k}^{p}$, any polynomial of degree $p$ can be represented by a single-hidden-layer $p$-th order MRePU network with at most $2p$ hidden nodes. $\square$
@@
Then, by replacing Lemma 40, which is used in the proof of Theorem 3 in (Shen23), we obtain the following theorem:
@@
Theorem 10 If $f:\mathbb{R}^{d}\rightarrow\mathbb{R}$ is a polynomial of $d$ variables with degree $N$, then $f$ can be reconstructed via MRePU neural network with depth $\mathcal{D}$, width $\mathcal{W}$, number of neurons $\mathcal{U}$ and total number of parameters $\mathcal{P}$ with following specification:
$\begin{split}&\mathcal{D}=2N-1,\quad\mathcal{W}=12pN^{d-1}+6p(N^{d-1}-N)/(N-1)\simeq\mathcal{O}(pN^{d-1}),\\ &\mathcal{U}=(6p+2)(2N^{d}-N^{d-1}-N)+2p(2N^{d}-N^{d-1}-N)/(N-1)\simeq\mathcal{O}(pN^{d}),\\ &\mathcal{P}=(30p+2)(2N^{d}-N^{d-1}-N)+(2p+1)(2N^{d}-N^{d-1}-N)/(N-1)\simeq\mathcal{O}(pN^{d}).\end{split}$
@@
Additionally, as a corollary to the above theorem, we obtain a universal approximation result for the class of differentiable functions, which corresponds to Theorem 5 in (Shen23). Before stating this result, we define a reference norm that will serve as our standard.
Definition 12. We define the norm of $f\in C^{s}(A)$, where $A\subset\mathbb{R}^{d}$, as following:
@@
Corollary 13. Let $f$ be a function belongs to $C^{s}(K)$ where $K\subset\mathbb{R}^{d}$ is a compact set. For any $N\in\mathbb{N}^{+}$, there exists a MRePU neural network $\psi_{N}$ with depth $\mathcal{D}$, width $\mathcal{W}$, number of neurons $\mathcal{U}$ and total number of parameters $\mathcal{P}$ with following specification:
@@ 5.2 Experiments on Synthetic Data
Regression on polynomials and differentiable functions First, we validate the efficiency of MRePU on polynomial functions. The neural network architecture consists of five hidden layers, each with a width of 32, and the input dimension is set according to the variables in each polynomial. For univariate polynomial problems, the training dataset was constructed with inputs from -2.0 to 2.0, divided into 101 points at intervals of 0.04. For $x^{3}-x$, the learning rate was set to 0.005, and $C_{W}$ was set to 0.5 for MRePU and 1.0 for other cases. For $x^{5}-5x^{3}+6x$, the learning rate was set to 0.0001, with $C_{W}$ set to 0.8 for MRePU case and learning rate to 0.001, $C_{W}$ to 1.0 for the others. For multivariate cases, the dataset for $xy$ was created by forming a meshgrid over the range $2,2]\times[-2,2]$, with each axis divided into 101 points. Similarly, for xyzxyz, a meshgrid was formed over the range $1,1]\times[-1,1]\times[-1,1]$, with each axis divided into 51 points. Training was conducted up to 10,000 epochs for all cases. The learning rate for multivariate cases was set to 0.001, and $C_{W}$ was set to 0.55 for $xy$, 0.6 for xyzxyz, and 1.0 for the other cases. Each hyperparameter was empirically determined to achieve optimal performance.
$x^{3}-x$ $x^{5}-5x^{3}+6x$ xyxy xyzxyz
MRePU $\mathbf{4.0\times 10^{-4}}$ $\mathbf{2.0\times 10^{-4}}$ $\mathbf{2.7\times 10^{-5}}$ $\mathbf{2.5\times 10^{-4}}$
ReLU $8.0\times 10^{-4}$ $7.0\times 10^{-4}$ $4.0\times 10^{-4}$ $4.2\times 10^{-2}$
GELU $7.2\times 10^{-2}$ $2.1\times 10^{-3}$ $4.0\times 10^{-4}$ $4.2\times 10^{-2}$
@@
Table 4 presents the average final loss over 25 experiments, where neural networks were trained to regress each polynomial function. As observed in the table, MRePU-based neural networks consistently outperform the others in all cases. Notably, for approximating xyzxyz, other networks fail to learn effectively, while only the MRePU network achieves meaningful learning. The loss plots for $x^{5}-5x^{3}+6x$ and $xy$ are depicted in Fig. 9. Fig. 10 illustrates the regression results of $x^{5}-5x^{3}+6x$ using ReLU and MRePU. A key observation is that the trained MRePU network approximates not only the function itself but also its derivative, whereas the ReLU network produces a piecewise function with noticeable discontinuities. Similarly, Fig. 11 shows the contour graphs of trained networks with different activation functions for $xy$. The MRePU network demonstrates a highly accurate approximation, while the ReLU network exhibits significant instability in the contours, and GELU shows reduced accuracy.
Figure 9: The training loss plots for each neural network over epochs for the cases of $x^{5}-5x^{3}+6x$ (Left) and xyxy (Right).
Figure 10: Top: The results of training for regressing $x^{5}-5x^{3}+6x$, where the left plot shows the output of the MRePU with $p=2$ network, and the right plot shows the output of the ReLU network. Bottom: The comparison between the derivative of the trained networks and the actual derivative function, $5x^{4}-15x^{2}+6$.
Figure 11: Contour plots for the reference function and the trained neural networks. Top left: MRePU with $p=2$ network. Top right: Reference (xyxy). Bottom left: GELU network. Bottom right: ReLU network.
@@
The training dataset for univariate functions was constructed in the same manner as for the polynomials. For multivariate functions, a meshgrid was generated by dividing the domain $1,1]\times[-1,1]$ into 51 intervals along each axis. The learning rate and $C_{W}$ values for these experiments were as follows: For $\lfloor x\rfloor_{+}^{3}$, the MRePU network used a learning rate of 0.0004 and $C_{W}=0.6$, while the other networks used 0.001 and $C_{W}=1.0$. For $f(x)$, The MRePU network employed a learning rate of 0.001 and $C_{W}=0.6$, while the other networks used 0.001 and $C_{W}=1.0$. For $g(x,y)$, the MRePU network used a learning rate of 0.002 and $C_{W}=0.8$, whereas the remaining networks used 0.003 and $C_{W}=1.0$. Each hyperparameter was empirically determined to achieve optimal performance.
@@
MRePU $8.1\times 10^{-5}$ $7.9\times 10^{-5}$ $4.0\times 10^{-4}$
ReLU $\mathbf{3.2\times 10^{-5}}$ $\mathbf{5.6\times 10^{-5}}$ $\mathbf{2.0\times 10^{-4}}$
GELU $2.4\times 10^{-3}$ $8.0\times 10^{-4}$ $4.4\times 10^{-1}$
@@
As shown in Table 5 and Fig. 12, for differentiable functions with discontinuities in higher-order derivatives, the ReLU network generally demonstrates the best performance. However, as observed in Figs. 13 and 14, even though the training loss of the ReLU network is lower (i.e., its $L^{2}$ norm is smaller), the MRePU network exhibits more stable function approximation in terms of contour smoothness and the behavior of the differentiated function. For the GELU network, it is evident that learning fails entirely for $g(x,y)$.
@@
Case of Physics-Informed Neural Networks (PINNs) Next, we compare and contrast neural networks constructed with different activation functions using the physics-informed neural networks (PINNs) (Raissi19) technique, which solves PDE problems by incorporating the PDE itself into the loss function. The PDE used in this experiment is the Burgers equation, given as:
@@
We defined the problem domain as $(x,t)\in[-1,1]\times[0,1]$ with $\nu=\frac{0.01}{\pi}$. The initial condition (the function value at $t=0$) was set to $\sin{\pi x}$, and the boundary condition was defined as u(1,t)=u(1,t)=0u(-1,t)=u(1,t)=0. The PINN loss comprises the $L^{2}$ loss for the boundary condition and the loss for the PDE. The number of points used for loss calculations was as follows: 10,000 randomly sampled collocation points within the domain and 100 points for the boundary condition. The hyperparameter settings included the use of the Adam optimizer with a learning rate of 0.001, betas set to $(0.9,0.999)$, and weight decay set to 0. For $C_{W}$, the value was set to 0.45 for MRePU with $p=2$ and 1.0 for all other cases. The neural network architecture was consistent across experiments, with an input dimension of 2, seven hidden layers, and a width of 64 per layer.
@@
The results of the experiment are presented in Fig. 15, 16, and 17. Interestingly, unlike the previous cases with polynomials and differentiable functions, the GELU activation exhibited the best performance, followed by the MRePU network. The ReLU network, however, failed to achieve meaningful learning. Additional experiments with the Tanh activation showed a performance similar to that of the GELU network. From this experiment, it can be concluded that the MRePU network is capable of achieving meaningful learning even under the highly complex training dynamics of the PINN loss (Wang22, Aditi21) and with a deep architecture consisting of 7 hidden layers (9 layers in total).
@@
Polynomial: MRePU >> ReLU >> GELU
Differentiable function: ReLU >> MRePU∗∗ >> GELU
PINN for Burgers equation: GELU >> MRePU >> ReLU
@@
5.3 Experiments on Real-World Data
In this section, we extend our experiments beyond synthetic data to real-world datasets, specifically MNIST and CIFAR-10. Consistent with the setup in Section 5.2, we trained neural network architectures employing MRePU, ReLU, and GELU activation functions. A particularly notable observation in this section is that MRePU networks can be directly applied to very deep neural networks and the widely used ResNet architecture with only minor modifications.
MNIST Classification First, we conducted experiments on the MNIST dataset, which is fundamental and widely used for verification. The $28\times 28$ input images were flattened into 784-dimensional vectors, and a 4-layer FCN architecture with dimensions $[784,256,256,$ $256,10]$ was employed. In all experiments, training was performed using the Adam optimizer with a learning rate of $10^{-3}$. For the first experiment verifying the criticality condition, a batch size of 512 was used, while in the second experiment comparing the performance of various activation functions, a batch size of 64 was utilized. In the first experiment, to empirically demonstrate that the condition in Eq. (11) serves as a boundary condition for training stability, the dataset was centered and scaled by dividing by $N_{K}$ (ranging from 0.1 to 2.0 with an interval of 0.1). Similarly, the hyperparameter $C_{W}$ was varied from 0.1 to 2.0 with an interval of 0.1. We performed training for 5 epochs for each case, repeated the process 10 times, and visualized the average test accuracy as a heatmap (Fig. 18).By regarding $C_{W}$ in Eq. (11) as a function $\tilde{C}_{W}(K)$ of $K$ for the case where $p=2$, the criticality condition or phase boundary obtained via fitting, $2.2\tilde{C}_{W}(0.02/N_{K})$, is depicted as a dotted red line in Fig. 18. As shown in the figure, the success of neural network training is highly sensitive to $C_{W}$, and the fitted threshold condition aligns well with the empirical observations. Notably, as $N_{K}$ increases and the data approaches the fixed point $K^{\star}=0$, the sensitivity to $C_{W}$ diminishes.
In the second experiment, for each activation function, the dataset was centered and normalized by dividing by $N_{K}=30$ to ensure proximity to the fixed point. The training process was conducted for 10 epochs and repeated 25 times. The average test loss and accuracy are presented in Table 6, while the training loss, plotted on a logarithmic scale with a 50-iteration moving average, is displayed in the left panel of Fig. 19. As illustrated in Fig. 19, MRePU demonstrates rapid convergence in the initial phase, followed by a convergence rate comparable to that of ReLU. Furthermore, the results in Table 6 indicate that all tested activation functions exhibit similar performance.
Figure 18: Phase diagram illustrating the criticality condition of the hyperparameter $C_{W}$ with respect to the normalization factor $N_{K}$. The background heatmap represents the empirical mean accuracy obtained from $10$ training results, where purple indicates higher performance. The overlaid red solid line denotes the theoretical phase boundary derived from Proposition 8, demonstrating a strong alignment with the empirical transition region.
CIFAR-10 Classification For the CIFAR-10 dataset, we employed the ResNet architecture. For ReLU and GELU, we adhered to the architecture described in (He15b), utilizing the basic building block. The network is constructed by arranging blocks with widths of 64, 128, 256, and 512 in a sequence of 2, 2, 1, and 1 blocks, respectively. Each block consists of two $3\times 3$ convolutional layers with a residual connection added to the input and output. Including the input and output layers, the total depth of the architecture is $1+2(2+2+1+1)=14$ layers.
For MRePU, we modified the residual structure as follows, which is an essential adjustment to maintain the kernel at a constant scale (Roberts22):
$z^{(l+1)}=\gamma\text{Block}(z^{(l)})+\sqrt{1-\gamma^{2}}z^{(l)}$
where $z^{(l+1)}$ is final output of whole block and $z^{(l)}$ is input of the block $0<\gamma<1$ is a weight for the residual term. Since MRePU, ReLU, and GELU all share $K^{\star}=0$ as a fixed point, it is fair to ensure the kernel scale remains close to zero. Consequently, we normalized the dataset to the following scale:
$X^{*}=\frac{X-\mathbb{E}[X]}{5000}$
We utilized a batch size of 64 and the Adam optimizer with a learning rate of $2.5\times 10^{-4}$. The training was conducted for 5 epochs. We performed a total of 25 independent trials for each activation function, and the average test loss and accuracy are reported in Table 6. The moving average of the log-scaled training loss, calculated over a window of 50 iterations, is depicted in the right panel of Fig. 19. While all three activation functions exhibit similar trends, it is observed that, unlike in the MNIST case, the reduction in training loss for MRePU lags slightly behind the others.
MNIST CIFAR-10
Loss Accuracy Loss Accuracy
MRePU $1.8\times 10^{-2}$ $97.7\%$ $1.29$ $57.7\%$
ReLU $\mathbf{1.3\times 10^{-2}}$ $\mathbf{98.4}\%$ $\mathbf{7.53}\times 10^{-1}$ $\mathbf{72.4}\%$
GELU $2.0\times 10^{-2}$ $98.0\%$ $7.71\times 10^{-1}$ $67.4\%$
Table 6: Mean final training losses and accuracies over 25 trials for real-world tasks for each of the MRePU (order $p=2$), ReLU, and GELU networks. The values in bold represent the lowest test loss values(highest for accuracies) in each column.
Figure 19: The training loss plots for each neural network over epochs for the cases of MNIST (Left) and CIFAR-10 (Right) datasets.